OpenStudy (tanabugg):

Please help!! write a function rule to represent the sequence 0.3,0.9,1.5,2.1... a. A(n)=0.6+(n-1)0.3 b. A(n)=0.6-(n)0.3 c. A(n)=0.3+(n-1)0.6 d. A(n)=0.3+(n+1)0.6

OpenStudy (nintendo):

Just test out the equations and see what works. Answer (a) is incorrect because \[0.6+(1-1)0.3=0.6 \neq 0.3\] Answer (b) is incorrect because for n=2, \[0.6-2 \cdot 0.3 = 0 \neq 0.9\] Answer (d) is incorrect because for n=1, \[0.3 + (1+1)0.6 = 1.5 \neq 0.3\] Therefore, the only valid option is (c).

OpenStudy (tanabugg):

write a function rule to represent the sequence 47,32,17,2 a. A(n)= 47 +(n - 1)15 b. A(n)= 47 +(n - 1)(-13) c. A(n)= 47+(n - 1)(-15) d. A(n)= 32 + (n - 1)15

OpenStudy (tanabugg):

That was the last one